What is the density (g/L) of H2 gas at a temperature of 283 K and 0.500 atm? (Hint: One approach is to assume a sample size of 1.0 L.)

-0.0107 g/L

-0.0217 g/L

-0.0434 g/L

-93.7 g/L

-0.0899 g/L
Answer #1

apply ideal gas law

PV = nRT

n = mass/MW

PV = mRT/MW

m/V = PMW/(RT)

D = P*MW/(RT)

D = (0.5*2)/(0.082*283) = 0.0430923 g/L

then

d = 0.0430923 which is answer C